2.) The EBITDA of Microsoft for the fiscal year ending Jun. 30, 2017 is closest to ($ in millions):

16,568

19,320

23,972 (Maybe)

31,104

(5.) Microsoft’s inventory turnover for the fiscal year ending Jun. 30, 2017 is closest to:

11.88

14.56

15.71 (Maybe)

18.67

(6.) Microsoft’s accounts receivable days for the fiscal year ending Jun. 30, 2017 is closest to:

65

80 (Maybe)

91

96

(8.) Microsoft’s asset turnover for the fiscal year ending Jun. 30, 2016 is closest to:

0.37

0.44 (Maybe)

0.53

0.66

(9.) Microsoft’s equity multiplier for the fiscal year ending Jun. 30, 2016 is closest to:

1.37

2.20

2.69 (Maybe)

3.33
Answer #1
2.) The EBITDA of Microsoft for the fiscal year ending Jun. 30, 2017 is closest to ($ in millions):
16,568
19,320
23,972
31,104
Microsoft Corp., EBITDA calculation
USD $ in millions
Net income $                                                               21,204.00
Add: Income tax expense $                                                                 1,945.00
Earnings before tax (EBT) $                                                               23,149.00
Add: Interest expense $                                                                 2,222.00
Earnings before interest and tax (EBIT) $                                                               25,371.00
Add: Depreciation, amortization, and other $                                                                 8,778.00
Earnings before interest, tax, depreciation and amortization (EBITDA) $                                                               34,149.00
(5.) Microsoft’s inventory turnover for the fiscal year ending Jun. 30, 2017 is closest to:
11.88
14.56
15.71
18.67
Inventory Turnover = COGS / Average Inventory
Inventory Turnover = 34,261 / ((2181+2251)/2) 15.46 times
(6.) Microsoft’s accounts receivable days for the fiscal year ending Jun. 30, 2017 is closest to:
65
80
91
96
Days Sales Outstanding= average Accounts Receivable/Revenue*Days in Period
Days Sales Outstanding = ((19,792 + 18,277)/2)/89950 x 365 80.31 days
(8.) Microsoft’s asset turnover for the fiscal year ending Jun. 30, 2016 is closest to:
0.37
0.44
0.53
0.66
Asset Turnover = Sales/Average Total Assets
Asset Turnover = 85320/193468 0.441
(9.) Microsoft’s equity multiplier for the fiscal year ending Jun. 30, 2016 is closest to:
1.37
2.2
2.69
3.33
Equity Multiplier = Total Assets/Total Shareholder equity
Equity Multiplier = 193468/71997 2.69